OpenStudy (anonymous):

Triangle GHM is isosceles. G (-2.-3), H (-2,7) and the x-coordinate of M is 4.? Find the five possible values of the -Y coordinate of M. GM=GH, GH=HM.

OpenStudy (anonymous):

Let the y-coordinate of M is y. Then applying in first case (i) GM = GH then (GM)^2 = (GH)^2 ((-2-4)^2)+((-3-y)^2) = ((-2+2)^2)+((-3-7)^2) 6^2 + (y+3)^2 = 0 + 10^2 (y+3)^2 = 100-36=64=8^2 y+3 = 8 or y+3 = -8 y = 5 or y= -11 Second Case (ii) GH = HM (GH)^2 = (HM)^2 ((-2+2)^2)+((-3-7)^2) = ((-2-4)^2)+((7-y)^2) 0 + 10^2 = 6^2 + (7-y)^2 (7-y)^2= 100-36 = 64 = 8^2 7-y = 8 or 7-y = -8 y = -1 or y = 15 Third Case (iii) GM = HM (GM)^2 = (HM)^2 ((-2-4)^2)+((-3-y)^2) = ((-2-4)^2)+((7-y)^2) 36 + (y+3)^2 = 36 + (y-7)^2 y + 3 = y-7 or y + 3 = 7 - y 3 = -7 not possible 2y = 4 y = 2. Therefore y can have 5 values -11, -1, 2, 5, 15.